HI I NEED HELP WITH A QUESTION I DON;T KNOW TO REACH THE CONCLSUION I KNOW NOTHING ABOUT PROPORTIONS THE QUESTION IS
SOLVE THE X IN THE FOLLOWING 3:80::X:150.

Answer by Alan3354(69443)

3:80::X:150
-----------
3/80 = x/150
Cross multiply
80x = 450
x = 45/8
-------------
You can also multiply and equate "means and extremes" in the ratio. Means are the 2 inner terms, extremes are the 2 outer.
80*x = 3*150 same thing